OpenStudy (anonymous):

Using Newton's Method, determine the first approximation to the solution of: e^(2.5x)=4−3x starting with x=0

OpenStudy (dumbcow):

if the function is e^2.5x +3x - 4 = 0 then the derivative is 2.5e^2.5x + 3 x1 = 0 - (e^2.5(0) + 3(0) -4)/(2.5e^2.5(0) +3) x1= 0 - (1-4)/(2.5+3) x1 = 3/5.5